OpenStudy (anonymous):

Find the geometric mean of 5 and 320. A. 35 B. 40 C. 45 D. 50

hartnn (hartnn):

Geometric Mean of a and b is \(\Large \sqrt {ab}\) so what will be Geometric mean of 5 and 320 ?

OpenStudy (anonymous):

ugh i m not really sure

hartnn (hartnn):

it'll be just \(\Large \sqrt{5\times 320} = ... ?\)

OpenStudy (imstuck):

Heres the set up for geometric mean, which we will call x:

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

715.54?

OpenStudy (imstuck):

\[\frac{ 5 }{ x }=\frac{ x }{ 120 }\]cross multiply and solve for x, your geometric mean.

hartnn (hartnn):

no, whats 320 * 5 = ... ?

OpenStudy (anonymous):

1600

hartnn (hartnn):

yes and its square root ?

OpenStudy (anonymous):

40

OpenStudy (anonymous):

thanks mate

OpenStudy (imstuck):

oops, I did 120...sorry! Set up is\[\frac{ 5 }{ x }=\frac{ x }{ 320 }\]

hartnn (hartnn):

\( \huge \checkmark \)

hartnn (hartnn):

welcome ^_^